OpenStudy (anonymous):

PLEASE SOLVE!! 4/5t + 2/5 = 3/5 My teacher is on skype with me and i suck at math however answering this will be the final question on moving on to my other grade! PLEASE ANSWER IT!! IT would mean loads, thanks <3

OpenStudy (anonymous):

4/5t + 2/5 = 3/5 Therefore 4t + 2=3 4t=3-2 t = 1/4